# The Map Equation Goes Neural: Mapping Network Flows with Graph Neural Networks

### v3

@@ B.3 Tabulated number of detected communities on real-world networks
Here, we report the average number of detected communities for Neuromap, DMoN, NOCD, DiffPool, MinCut, Ortho, and Infomap on real-world networks. Table 4 lists the results on the Cora, CiteSeer, PubMed, Amazon Computer (PC), and Amazon Photo datasets. Table 5 lists the results on the Coauthor CS, Coatuhor Physics, Cora ML, Wiki CS, and ogb-arxiv datasets.
@@
Table 4: Average number of detected communities and their standard deviations on real-world networks. We tested each method with 2-layer MLP, GCN, GIN, and SAGE architectures, except for Infomap, which is not based on deep learning.
@@
Method Arch. Cora CiteSeer PubMed PC Photo
Neuromap LIN $14.4\pm 3.7$ $38.0\pm 3.9$ $12.4\pm 8.1$ $1.0\pm 0.0$ $1.0\pm 0.0$
MLP $44.4\pm 2.5$ $57.3\pm 0.9$ $30.8\pm 3.8$ $3.0\pm 0.8$ $4.0\pm 1.2$
GCN $50.7\pm 0.9$ $58.0\pm 0.0$ $51.6\pm 4.4$ $21.2\pm 3.4$ $19.6\pm 2.1$
GIN $20.8\pm 2.2$ $32.0\pm 2.7$ $12.7\pm 1.3$ $6.4\pm 1.3$ $6.8\pm 0.8$
SAGE $43.9\pm 2.4$ $57.6\pm 0.6$ $34.4\pm 2.5$ $7.2\pm 1.9$ $8.1\pm 1.3$
DMoN LIN $34.6\pm 8.0$ $24.1\pm 9.4$ $2.3\pm 1.6$ $3.4\pm 1.3$ $3.6\pm 1.6$
MLP $52.0\pm 0.0$ $58.0\pm 0.0$ $125.7\pm 3.3$ $78.9\pm 6.7$ $78.0\pm 3.0$
GCN $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $116.8\pm 0.4$ $86.9\pm 0.3$
GIN $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $114.1\pm 1.8$ $86.8\pm 0.4$
SAGE $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $116.8\pm 0.4$ $86.6\pm 0.7$
NOCD LIN $20.5\pm 10.2$ $17.5\pm 21.5$ $12.1\pm 3.1$ $1.7\pm 1.0$ $1.0\pm 0.2$
MLP $29.3\pm 2.7$ $23.6\pm 4.1$ $29.8\pm 2.7$ $13.9\pm 3.9$ $11.0\pm 2.5$
GCN $51.7\pm 0.6$ $58.0\pm 0.0$ $35.1\pm 2.9$ $72.3\pm 19.2$ $51.1\pm 11.0$
GIN $24.2\pm 1.7$ $28.8\pm 2.6$ $35.9\pm 6.2$ $33.1\pm 5.8$ $25.1\pm 4.0$
SAGE $47.8\pm 1.6$ $56.6\pm 1.3$ $60.6\pm 3.3$ $32.0\pm 3.0$ $25.1\pm 3.0$
DiffPool LIN $2.6\pm 1.1$ $4.6\pm 3.3$ $1.0\pm 0.0$ $1.2\pm 0.5$ $1.0\pm 0.0$
MLP $51.8\pm 0.4$ $58.0\pm 0.2$ $43.6\pm 4.7$ $2.2\pm 0.5$ $2.2\pm 0.4$
GCN $48.2\pm 1.6$ $57.9\pm 0.4$ $15.6\pm 2.1$ $5.7\pm 1.1$ $5.7\pm 0.8$
GIN $11.5\pm 2.2$ $17.4\pm 3.1$ $7.3\pm 1.6$ $3.5\pm 0.6$ $5.2\pm 0.8$
SAGE $49.7\pm 1.3$ $58.0\pm 0.2$ $26.0\pm 3.0$ $3.2\pm 0.4$ $3.9\pm 0.7$
MinCut LIN $51.3\pm 0.9$ $54.8\pm 1.9$ $40.0\pm 6.1$ $1.0\pm 0.2$ $1.1\pm 0.3$
MLP $51.5\pm 1.6$ $58.0\pm 0.0$ $25.7\pm 7.4$ $1.0\pm 0.0$ $1.2\pm 0.4$
GCN $52.0\pm 0.0$ $58.0\pm 0.0$ $134.9\pm 4.1$ $69.5\pm 25.2$ $62.8\pm 20.3$
GIN $21.5\pm 20.0$ $55.7\pm 8.5$ $15.0\pm 37.5$ $11.0\pm 4.2$ $10.1\pm 5.5$
SAGE $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $115.9\pm 2.8$ $86.9\pm 0.3$
Ortho LIN $48.7\pm 1.3$ $53.4\pm 1.7$ $68.1\pm 5.7$ $9.5\pm 4.1$ $9.4\pm 5.1$
MLP $52.0\pm 0.0$ $58.0\pm 0.0$ $126.9\pm 4.7$ $8.8\pm 2.3$ $11.2\pm 2.0$
GCN $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $114.9\pm 1.6$ $85.0\pm 3.3$
GIN $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $95.8\pm 4.4$ $78.8\pm 2.8$
SAGE $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $52.5\pm 30.2$ $52.4\pm 21.7$
Infomap $282.2\pm 4.1$ $628.9\pm 2.7$ $924.9\pm 9.1$ $455.1\pm 5.1$ $223.5\pm 4.4$
Table 5: Average number of detected communities and their standard deviations on real-world networks. We tested each method with 2-layer MLP, GCN, GIN, and SAGE architectures, except for Infomap, which is not based on deep learning.
Method Arch. CS Physics Cora ML Wiki CS ogb-arxiv
Neuromap LIN $50.1\pm 5.2$ $34.5\pm 4.5$ $49.0\pm 3.2$ $15.1\pm 1.9$ $3.9\pm 2.6$
MLP $22.2\pm 2.4$ $35.8\pm 3.3$ $54.4\pm 0.7$ $14.2\pm 2.0$ $29.1\pm 5.4$
GCN $34.6\pm 4.5$ $73.0\pm 8.0$ $55.0\pm 0.0$ $27.4\pm 2.9$ $34.8\pm 5.4$
GIN $14.2\pm 2.6$ $14.9\pm 1.8$ $43.3\pm 3.5$ $8.6\pm 1.8$ $5.1\pm 0.8$
SAGE $22.0\pm 2.8$ $41.3\pm 6.1$ $54.9\pm 0.3$ $78.0\pm 6.9$ $38.3\pm 3.5$
DMoN LIN $14.4\pm 5.9$ $10.2\pm 4.5$ $15.0\pm 18.6$ $11.0\pm 5.8$ $10.1\pm 9.9$
MLP $120.2\pm 4.1$ $141.7\pm 7.4$ $55.0\pm 0.0$ $99.5\pm 3.5$ $166.6\pm 9.1$
GCN $135.0\pm 0.0$ $186.0\pm 0.0$ $55.0\pm 0.0$ $108.0\pm 0.0$ $400.1\pm 3.1$
GIN $134.7\pm 0.4$ $185.4\pm 0.6$ $55.0\pm 0.0$ $106.6\pm 1.5$ $278.2\pm 18.2$
SAGE $135.0\pm 0.0$ $186.0\pm 0.2$ $55.0\pm 0.0$ $108.0\pm 0.0$ $409.2\pm 2.1$
NOCD LIN $34.7\pm 8.5$ $21.6\pm 2.4$ $24.8\pm 3.9$ $11.2\pm 0.9$ $12.9\pm 2.8$
MLP $20.4\pm 3.7$ $22.8\pm 2.7$ $27.6\pm 1.7$ $28.3\pm 2.0$ $145.4\pm 41.7$
GCN $68.4\pm 8.6$ $73.9\pm 7.3$ $54.0\pm 0.8$ $33.8\pm 3.9$ $69.0\pm 6.5$
GIN $27.0\pm 4.1$ $29.9\pm 4.3$ $24.9\pm 2.6$ $45.7\pm 5.1$ OOM
SAGE $45.3\pm 3.6$ $55.7\pm 2.9$ $51.1\pm 1.7$ $46.6\pm 2.7$ $68.0\pm 6.6$
DiffPool LIN $6.7\pm 6.1$ $8.1\pm 5.5$ $1.0\pm 0.2$ $4.4\pm 1.0$ OOM
MLP $16.5\pm 2.6$ $39.8\pm 6.1$ $55.0\pm 0.0$ $10.0\pm 1.4$ OOM
GCN $12.8\pm 2.3$ $19.0\pm 3.3$ $54.1\pm 0.9$ $7.7\pm 1.4$ OOM
GIN $7.0\pm 1.1$ $6.6\pm 1.1$ $15.5\pm 2.5$ $4.7\pm 1.0$ OOM
SAGE $10.7\pm 1.6$ $15.8\pm 2.1$ $54.2\pm 0.7$ $10.1\pm 1.2$ OOM
MinCut LIN $91.0\pm 3.9$ $89.6\pm 6.2$ $55.0\pm 0.0$ $48.5\pm 4.9$ $81.6\pm 93.0$
MLP $21.2\pm 41.8$ $118.6\pm 24.6$ $38.1\pm 5.8$ $101.4\pm 19.2$ $155.0\pm 49.5$
GCN $133.2\pm 2.9$ $157.0\pm 40.2$ $55.0\pm 0.0$ $108.0\pm 0.0$ $97.4\pm 20.9$
GIN $73.9\pm 63.8$ $6.1\pm 6.7$ $40.6\pm 6.0$ $4.0\pm 14.5$ $10.5\pm 4.2$
SAGE $135.0\pm 0.0$ $176.5\pm 24.6$ $55.0\pm 0.0$ $108.0\pm 0.0$ $379.5\pm 33.7$
Ortho LIN $97.4\pm 6.6$ $149.5\pm 5.4$ $34.5\pm 3.3$ $33.2\pm 2.6$ $32.8\pm 3.4$
MLP $48.1\pm 9.1$ $101.7\pm 11.2$ $55.0\pm 0.0$ $41.0\pm 6.2$ $187.7\pm 18.9$
GCN $115.6\pm 11.9$ $166.2\pm 15.9$ $55.0\pm 0.0$ $108.0\pm 0.0$ $253.8\pm 68.3$
GIN $135.0\pm 0.0$ $185.8\pm 1.0$ $55.0\pm 0.0$ $107.5\pm 0.8$ $170.2\pm 22.4$
SAGE $84.9\pm 17.5$ $109.5\pm 18.1$ $55.0\pm 0.0$ $105.1\pm 1.6$ $379.8\pm 4.6$
Infomap $834.4\pm 9.4$ $1160.8\pm 11.4$ $287.4\pm 3.8$ $747.0\pm 8.4$ $4840.0\pm 22.4$
Appendix C Results on real-world networks with fixed number of hidden channels
Here we provide further results on real-world networks with a different configuration of the neural networks: we set the number of hidden channels to 512 and set the maximum number of communities to the “ground-truth” number of communities, that is $s=\left|Y\right|$.
C.1 Average Achieved AMI
@@
C.2 Average Number of Detected Communities
@@
C.3 Tabulated AMI results on real-world networks
Table 6: Average AMI in % (higher is better) and their standard deviations on real-world networks. For each dataset, we highlight the best scores in bold and underline the second-best score. We tested each method with 2-layer MLP, GCN, GIN, and SAGE architectures, except for Infomap, which is not based on deep learning.
Method Arch. Cora CiteSeer PubMed PC Photo
Neuromap LIN $19.9\pm 10.3$ $12.8\pm 6.9$ $0.9\pm 3.7$ $-0.0\pm 0.0$ $0.0\pm 0$ 0.0
MLP $32.7\pm 4.6$ $13.6\pm 2.6$ $\textbf{22.9}\pm 4.6$ $33.6\pm 13.3$ $51.3\pm 11.1$
GCN $31.7\pm 3.0$ $12.8\pm 2.9$ $19.2\pm 5.8$ $\textbf{49.9}\pm 3.6$ $61.4\pm 4.2$
GIN $\textbf{35.8}\pm 4.1$ $16.1\pm 3.2$ $16.1\pm 5.8$ $40.5\pm 3.9$ $50.3\pm 7.0$
SAGE $32.5\pm 5.0$ $12.9\pm 2.7$ $19.8\pm 4.6$ $45.5\pm 5.4$ $49.2\pm 6.1$
DMoN LIN $2.4\pm 1.4$ $1.9\pm 1.3$ $1.3\pm 2.0$ $4.3\pm 2.5$ $7.1\pm 4.2$
MLP $8.9\pm 2.4$ $6.7\pm 2.5$ $17.0\pm 5.7$ $14.9\pm 5.4$ $17.4\pm 5.8$
GCN $30.4\pm 4.3$ $16.8\pm 3.5$ $17.3\pm 6.3$ $43.3\pm 2.7$ $54.0\pm 3.7$
GIN $15.0\pm 4.8$ $8.6\pm 3.3$ $8.2\pm 5.2$ $27.1\pm 2.5$ $31.1\pm 5.6$
SAGE $30.3\pm 3.9$ $16.0\pm 2.9$ $15.9\pm 4.1$ $29.7\pm 7.0$ $37.9\pm 10.1$
NOCD LIN $2.9\pm 1.3$ $3.0\pm 1.6$ $0.2\pm 0.5$ $1.9\pm 2.5$ $2.1\pm 3.0$
MLP $33.1\pm 14.7$ $\textbf{29.6}\pm 10.8$ $16.8\pm 8.1$ $47.9\pm 2.1$ $55.1\pm 15.4$
GCN $32.6\pm 4.6$ $11.7\pm 1.7$ $19.3\pm 7.3$ $\underline{49.6}\pm 2.2$ $\textbf{64.4}\pm 4.1$
GIN $35.1\pm 4.8$ $19.2\pm 3.4$ $17.0\pm 6.3$ $32.6\pm 3.7$ $43.3\pm 4.1$
SAGE $32.5\pm 3.7$ $11.6\pm 2.8$ $\underline{22.6}\pm 4.9$ $47.7\pm 2.4$ $\underline{63.3}\pm 3.5$
DiffPool LIN $0.8\pm 1.9$ $1.2\pm 0.5$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
MLP $4.9\pm 1.3$ $4.1\pm 1.4$ $10.3\pm 3.6$ $1.6\pm 0.4$ $3.5\pm 1.4$
GCN $28.5\pm 4.1$ $12.4\pm 2.5$ $18.8\pm 5.5$ $30.4\pm 9.1$ $46.4\pm 6.5$
GIN $14.0\pm 4.8$ $8.2\pm 3.1$ $2.1\pm 1.7$ $5.3\pm 5.6$ $5.7\pm 6.8$
SAGE $24.4\pm 3.0$ $11.6\pm 2.5$ $18.9\pm 5.0$ $3.1\pm 0.6$ $12.8\pm 6.1$
MinCut LIN $30.6\pm 4.3$ $16.7\pm 2.6$ $21.0\pm 1.9$ $3.1\pm 8.5$ $1.9\pm 6.4$
MLP $30.4\pm 4.5$ $15.5\pm 3.5$ $19.7\pm 6.3$ $38.8\pm 13.3$ $56.4\pm 2.7$
GCN $31.9\pm 3.7$ $13.1\pm 2.8$ $18.5\pm 4.2$ $46.3\pm 2.0$ $56.1\pm 2.9$
GIN $\underline{35.7}\pm 2.9$ $18.9\pm 3.0$ $20.2\pm 4.1$ $32.4\pm 7.8$ $49.2\pm 4.9$
SAGE $33.6\pm 2.9$ $13.9\pm 3.5$ $17.1\pm 3.7$ $43.7\pm 2.5$ $54.0\pm 2.8$
Ortho LIN $8.2\pm 1.2$ $8.0\pm 2.3$ $15.7\pm 4.6$ $12.4\pm 6.5$ $13.6\pm 8.6$
MLP $4.4\pm 1.0$ $3.5\pm 1.0$ $8.5\pm 3.6$ $22.4\pm 5.7$ $27.1\pm 2.9$
GCN $22.0\pm 3.0$ $10.2\pm 2.1$ $16.4\pm 4.8$ $45.1\pm 1.9$ $54.4\pm 3.6$
GIN $19.2\pm 3.6$ $9.4\pm 2.9$ $6.3\pm 3.0$ $37.4\pm 3.8$ $45.9\pm 3.0$
SAGE $21.5\pm 4.1$ $7.6\pm 2.5$ $15.9\pm 3.8$ $29.4\pm 2.3$ $36.6\pm 3.2$
Infomap $35.2\pm 0.2$ $\underline{23.6}\pm 0.1$ $16.0\pm 0.1$ $49.5\pm 0.5$ $57.3\pm 1.1$

### v4

@@ C.5 Tabulated number of detected communities on real-world networks
Here, we report the average number of detected communities for Neuromap, DMoN, NOCD, DiffPool, MinCut, Ortho, and Infomap on real-world networks for setting the number of hidden dimensions to $4\sqrt{n}$ and the maximum number of communities to $s=\sqrt{n}$.
@@
Table 5: Average number of detected communities and their standard deviations on real-world networks. We tested each method with 2-layer MLP, GCN, GIN, and SAGE architectures, except for Infomap, which is not based on deep learning. OOM stands for “out of memory”.
@@
Method Arch. Cora CiteSeer PubMed PC Photo CS Physics Cora ML Wiki CS ogb-arxiv
Neuromap LIN $14.4\pm 3.7$ $38.0\pm 3.9$ $12.4\pm 8.1$ $1.0\pm 0.0$ $1.0\pm 0.0$ $50.1\pm 5.2$ $34.5\pm 4.5$ $11.4\pm 2.5$ $17.0\pm 1.7$ $2.4\pm 1.6$
MLP $44.4\pm 2.5$ $57.3\pm 0.9$ $30.8\pm 3.8$ $3.0\pm 0.8$ $4.0\pm 1.2$ $22.2\pm 2.4$ $35.8\pm 3.3$ $48.3\pm 1.9$ $12.3\pm 1.3$ $29.5\pm 4.4$
GCN $50.7\pm 0.9$ $58.0\pm 0.0$ $51.6\pm 4.4$ $21.2\pm 3.4$ $19.6\pm 2.1$ $34.6\pm 4.5$ $73.0\pm 8.0$ $54.3\pm 0.7$ $17.5\pm 2.4$ $34.4\pm 7.0$
GIN $20.8\pm 2.2$ $32.0\pm 2.7$ $12.7\pm 1.3$ $6.4\pm 1.3$ $6.8\pm 0.8$ $14.2\pm 2.6$ $14.9\pm 1.8$ $22.6\pm 2.7$ $6.8\pm 1.7$ $5.5\pm 1.3$
SAGE $43.9\pm 2.4$ $57.6\pm 0.6$ $34.4\pm 2.5$ $7.2\pm 1.9$ $8.1\pm 1.3$ $22.0\pm 2.8$ $41.3\pm 6.1$ $45.4\pm 2.8$ $17.1\pm 2.0$ $39.5\pm 4.4$
DMoN LIN $13.6\pm 13.3$ $25.1\pm 10.8$ $2.5\pm 1.7$ $2.4\pm 1.0$ $2.6\pm 1.4$ $10.0\pm 6.3$ $7.7\pm 6.7$ $4.5\pm 6.2$ $10.9\pm 5.6$ $7.6\pm 6.4$
MLP $52.0\pm 0.2$ $57.9\pm 0.3$ $24.8\pm 33.0$ $51.4\pm 36.3$ $81.2\pm 15.0$ $63.1\pm 17.1$ $76.3\pm 6.7$ $55.0\pm 0.2$ $92.9\pm 27.1$ $118.9\pm 13.2$
GCN $51.9\pm 0.3$ $58.0\pm 0.0$ $140.0\pm 0.0$ $117.0\pm 0.0$ $87.0\pm 0.0$ $135.0\pm 0.0$ $186.0\pm 0.0$ $54.8\pm 0.4$ $108.0\pm 0.0$ $412.0\pm 0.0$
GIN $51.6\pm 0.6$ $57.0\pm 1.0$ $140.0\pm 0.0$ $110.6\pm 6.1$ $84.5\pm 1.7$ $134.9\pm 0.3$ $185.9\pm 0.3$ $54.5\pm 0.6$ $97.2\pm 4.0$ $301.8\pm 77.3$
SAGE $51.8\pm 0.4$ $58.0\pm 0.0$ $140.0\pm 0.2$ $116.8\pm 0.7$ $86.8\pm 0.4$ $135.0\pm 0.0$ $186.0\pm 0.0$ $54.8\pm 0.4$ $108.0\pm 0.0$ $412.0\pm 0.0$
NOCD LIN $20.5\pm 10.2$ $17.5\pm 21.5$ $12.1\pm 3.1$ $1.7\pm 1.0$ $1.0\pm 0.2$ $34.7\pm 8.5$ $21.6\pm 2.4$ $24.8\pm 3.9$ $11.2\pm 0.9$ $12.9\pm 2.8$
MLP $29.3\pm 2.7$ $23.6\pm 4.1$ $29.8\pm 2.7$ $13.9\pm 3.9$ $11.0\pm 2.5$ $20.4\pm 3.7$ $22.8\pm 2.7$ $27.6\pm 1.7$ $28.3\pm 2.0$ $145.4\pm 41.7$
GCN $51.7\pm 0.6$ $58.0\pm 0.0$ $35.1\pm 2.9$ $72.3\pm 19.2$ $51.1\pm 11.0$ $68.4\pm 8.6$ $73.9\pm 7.3$ $54.0\pm 0.8$ $33.8\pm 3.9$ $69.0\pm 6.5$
GIN $24.2\pm 1.7$ $28.8\pm 2.6$ $35.9\pm 6.2$ $33.1\pm 5.8$ $25.1\pm 4.0$ $27.0\pm 4.1$ $29.9\pm 4.3$ $24.9\pm 2.6$ $45.7\pm 5.1$ $78.5\pm 15.6$
SAGE $47.8\pm 1.6$ $56.6\pm 1.3$ $60.6\pm 3.3$ $32.0\pm 3.0$ $25.1\pm 3.0$ $45.3\pm 3.6$ $55.7\pm 2.9$ $51.1\pm 1.7$ $46.6\pm 2.7$ $68.0\pm 6.6$
DiffPool LIN $2.6\pm 1.1$ $4.6\pm 3.3$ $1.0\pm 0.0$ $1.2\pm 0.5$ $1.0\pm 0.0$ $6.7\pm 6.1$ $8.1\pm 5.5$ $1.0\pm 0.2$ $4.4\pm 1.0$ OOM
MLP $51.8\pm 0.4$ $58.0\pm 0.2$ $43.6\pm 4.7$ $2.2\pm 0.5$ $2.2\pm 0.4$ $16.5\pm 2.6$ $39.8\pm 6.1$ $55.0\pm 0.0$ $10.0\pm 1.4$ OOM
GCN $48.2\pm 1.6$ $57.9\pm 0.4$ $15.6\pm 2.1$ $5.7\pm 1.1$ $5.7\pm 0.8$ $12.8\pm 2.3$ $19.0\pm 3.3$ $54.1\pm 0.9$ $7.7\pm 1.4$ OOM
GIN $11.5\pm 2.2$ $17.4\pm 3.1$ $7.3\pm 1.6$ $3.5\pm 0.6$ $5.2\pm 0.8$ $7.0\pm 1.1$ $6.6\pm 1.1$ $15.5\pm 2.5$ $4.7\pm 1.0$ OOM
SAGE $49.7\pm 1.3$ $58.0\pm 0.2$ $26.0\pm 3.0$ $3.2\pm 0.4$ $3.9\pm 0.7$ $10.7\pm 1.6$ $15.8\pm 2.1$ $54.2\pm 0.7$ $10.1\pm 1.2$ OOM
MinCut LIN $51.3\pm 0.9$ $54.8\pm 1.9$ $40.0\pm 6.1$ $1.0\pm 0.2$ $1.1\pm 0.3$ $91.0\pm 3.9$ $89.6\pm 6.2$ $55.0\pm 0.0$ $48.5\pm 4.9$ $81.6\pm 93.0$
MLP $51.5\pm 1.6$ $58.0\pm 0.0$ $25.7\pm 7.4$ $1.0\pm 0.0$ $1.2\pm 0.4$ $21.2\pm 41.8$ $118.6\pm 24.6$ $38.1\pm 5.8$ $101.4\pm 19.2$ $155.0\pm 49.5$
GCN $52.0\pm 0.0$ $58.0\pm 0.0$ $134.9\pm 4.1$ $69.5\pm 25.2$ $62.8\pm 20.3$ $133.2\pm 2.9$ $157.0\pm 40.2$ $55.0\pm 0.0$ $108.0\pm 0.0$ $97.4\pm 20.9$
GIN $21.5\pm 20.0$ $55.7\pm 8.5$ $15.0\pm 37.5$ $11.0\pm 4.2$ $10.1\pm 5.5$ $73.9\pm 63.8$ $6.1\pm 6.7$ $40.6\pm 6.0$ $4.0\pm 14.5$ $10.5\pm 4.2$
SAGE $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $115.9\pm 2.8$ $86.9\pm 0.3$ $135.0\pm 0.0$ $176.5\pm 24.6$ $55.0\pm 0.0$ $108.0\pm 0.0$ $379.5\pm 33.7$
Ortho LIN $48.7\pm 1.3$ $53.4\pm 1.7$ $68.1\pm 5.7$ $9.5\pm 4.1$ $9.4\pm 5.1$ $97.4\pm 6.6$ $149.5\pm 5.4$ $34.5\pm 3.3$ $33.2\pm 2.6$ $32.8\pm 3.4$
MLP $52.0\pm 0.0$ $58.0\pm 0.0$ $126.9\pm 4.7$ $8.8\pm 2.3$ $11.2\pm 2.0$ $48.1\pm 9.1$ $101.7\pm 11.2$ $55.0\pm 0.0$ $41.0\pm 6.2$ $187.7\pm 18.9$
GCN $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $114.9\pm 1.6$ $85.0\pm 3.3$ $115.6\pm 11.9$ $166.2\pm 15.9$ $55.0\pm 0.0$ $108.0\pm 0.0$ $253.8\pm 68.3$
GIN $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $95.8\pm 4.4$ $78.8\pm 2.8$ $135.0\pm 0.0$ $185.8\pm 1.0$ $55.0\pm 0.0$ $107.5\pm 0.8$ $170.2\pm 22.4$
SAGE $52.0\pm 0.0$ $58.0\pm 0.0$ $140.0\pm 0.0$ $52.5\pm 30.2$ $52.4\pm 21.7$ $84.9\pm 17.5$ $109.5\pm 18.1$ $55.0\pm 0.0$ $105.1\pm 1.6$ $379.8\pm 4.6$
Infomap $282.2\pm 4.1$ $628.9\pm 2.7$ $924.9\pm 9.1$ $455.1\pm 5.1$ $223.5\pm 4.4$ $834.4\pm 9.4$ $1160.8\pm 11.4$ $287.4\pm 3.8$ $747.0\pm 8.4$ $4840.0\pm 22.4$
Appendix D Results on real-world networks with fixed number of hidden channels
We provide further results on real-world networks with a different configuration of the neural networks: we set the number of hidden channels to 512 and set the maximum number of communities to the “ground-truth” number of communities, that is $s=\left|Y\right|$.
D.1 Average Achieved AMI
@@
D.2 Average Number of Detected Communities
@@
D.3 Tabulated AMI results on real-world networks
Here, we report the detailed AMI performance for Neuromap, DMoN, NOCD, DiffPool, MinCut, Ortho, and Infomap on real-world networks for setting the number of hidden dimensions to $512$ and the maximum number of communities to $s=\left|Y\right|$. When several of the best-performing methods achieve similar AMI, we use an independent two-sample t-test to determine whether one of them can be considered to perform better than the other. In cases where their performances do not differ significantly, we mark both as best.